It's almost Valentine's Day and Pastry Garden is gearing up. This year they are offering gift baskets at a cost of $125. The variable cost is $60. Pastry Garden pays salaries of $1,300 and the store rent is $1,000 per month. There is also a sales commission of 5% of the sales price. They expect to sell 140 gift baskets.

What is the total revenue from gift baskets?

What is the total cost of goods sold on the gift baskets?

What is the commission per unit?

What is the total commission?
Answer #1
1
Total revenue from gift baskets 17500 =140*125
2
Total cost of goods sold on the gift baskets 8400 =140*60
3
Commission per unit 6.25 =125*5%
4
Total commission 875 =140*6.25